OpenStudy (anonymous):

How would I simplify (3b^3/2a^3b^5)^4

OpenStudy (anonymous):

use this rule: \((a^n)^m=a^{nm}\)

OpenStudy (anonymous):

u tried @maddiemaze ??

OpenStudy (anonymous):

I'm not sure how to apply the rule

OpenStudy (anonymous):

\[\huge\frac{ 81a^{12}b^{32} }{ 16 }\]

OpenStudy (anonymous):

i do the nominator for you: \(81b^{12}\)

OpenStudy (anonymous):

\[\huge (\frac{ a }{ b })^m=\frac{ a^m }{ b^m }\]

OpenStudy (anonymous):

okay @myko if the numerator is 81b^12 would the denominator be 16b^32

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@deena123 would one of the letters be at the bottom?

OpenStudy (anonymous):

\(2^4(a^3)^4(b^5)^4=?\)

OpenStudy (anonymous):

sorry answer should be \[\huge (\frac{ 3b^3 }{ 2a^3b^5 })^4=\frac{ 81 }{ 16a^{12}b^8 }\]

OpenStudy (anonymous):

@myko ??

OpenStudy (anonymous):

that's the denominator

OpenStudy (anonymous):

that should be the answer!!

OpenStudy (anonymous):

so: \((3b^3/2a^3b^5)^4=\frac{81b^{12}}{16a^{12}b^{20}}=\frac{81}{16a^{12}b^8}\)

OpenStudy (anonymous):

Thanks guys

OpenStudy (anonymous):

yw